OpenStudy (anonymous):

Multiply and simplify by factoring.... √12 * √150

OpenStudy (anonymous):

2root 3 *5root 6 10root 18 20root 2 tats it simple

OpenStudy (vishweshshrimali5):

20 root 2

OpenStudy (anonymous):

20*1.4=28

OpenStudy (anonymous):

I do not understand how this is done?

OpenStudy (anonymous):

why is it 20sqrt 2?

OpenStudy (anonymous):

\[\sqrt{12}\times \sqrt{150}=\sqrt{1800}\rightarrow \sqrt{900\times2}=30\sqrt{2}\]

OpenStudy (anonymous):

Thanks Nalita.... that definitely makes more sense and its correct!!!!!!